# Watch Out Your Album! On the Inadvertent Privacy Memorization in Multi-Modal Large Language Models

### v1

@@ 1 Introduction
Figure 1: An overview for contrasting task-irrelevant private content (red) with task-relevant private content (blue and green) commonly examined in prior studies. The example image is sampled from the Android in the Wild dataset (Rawles et al., 2023). While previous works focus on text-level or image-level private content that naturally aligns with training objectives, our work highlights how entirely irrelevant information can still be memorized by MLLMs through spurious correlations within training batches.
@@ 2 Preliminary: Task-Irrelevant Content
Figure 2: An illustration of how MLLMs can inadvertently memorize private content that is globally irrelevant yet has a high probability of forming spurious correlations with downstream tasks in a mini-batch (Proof in Appendix A). Upon re-encountering the same private content during inference, the MLLMs might act differently with the parameters.
@@ 5.1 Privacy Concerns in LLMs
Recent research has sought to understand the extent to which LLMs memorize and potentially leak sensitive training data (Li et al., 2023a; Satvaty et al., 2024; Ippolito et al., 2023). A central line of research involves probing LLMs with carefully crafted prompts to expose memorized sequences that resemble personal identifiers or private user information (Lukas et al., 2023; Carlini et al., 2023). Carlini et al. (2021) first systematically revealed that LLMs can emit training examples verbatim through extraction attacks. Subsequently, Tirumala et al. (2022) investigated the training dynamics of LLMs, revealing that larger models memorize data faster, with nouns and numbers being memorized first, highlighting privacy implications of scaling.
@@ 5.2 Privacy Concerns in MLLMs
Compared to LLMs, privacy concerns in MLLMs remain less explored. Pinto et al. (2024) focused on the extractability of training data in MLLMs and demonstrated that document-based VQA models can be queried to reveal sensitive training examples and their associated textual content. Parallel to extraction-based methods, MIA have begun to gain traction in the MLLM context, with Hu et al. (2022b) providing an early attempt. Following this line, Ko et al. (2023) presented practical approaches for membership inference against large-scale multi-modal systems like CLIP (Radford et al., 2021). Recently, Li et al. (2024) introduced the first systematic benchmarking of MIA for large vision-language models (VLLMs), unveiling new challenges specific to the multi-modal domain.
@@ 6 Discussion and Future Directions
From the defender’s perspective, our paper suggests increasing batch sizes or using gradient accumulation to mitigate the inadvertent memorization of spurious correlations. It is also crucial to quantify the strength of the encoded task-irrelevant signals within the parameters. Future work could investigate the model-specific lower bound on safe batch sizes that sufficiently limit inadvertent task-irrelevant memorization.
@@ 7 Conclusion
In this paper, we investigate a critical yet underexplored question regarding whether MLLMs memorize private content that is entirely irrelevant to downstream tasks. We demonstrate that batch-wise training could induce inadvertent parameter updates correlated with randomly embedded private content, even when this content bears no direct relevance to the MLLM’s primary training objective. Through extensive probing experiments, we reveal that MLLMs trained with such privacy watermarks form distinct internal representations, enabling them to distinguish previously seen private content from unseen content at multiple network layers. Notably, we found that while MLLMs do not necessarily reproduce the memorized knowledge through direct prompting, they nonetheless encode these task-irrelevant details in their parameter space. Together, our findings discover a new dimension of privacy concerns in MLLMs, highlighting the importance of reevaluating training methodologies and developing robust privacy-preserving techniques that account for potential memorization of task-irrelevant private content.
@@ Appendix A Mathematical Proof of Spurious Correlations in Mini-Batch Training
where $\overline{\mathbf{y}}=\frac{1}{B}\sum_{i=1}^{B}\mathbf{y}_{i}$ and $\overline{\mathbf{u}}=\frac{1}{B}\sum_{i=1}^{B}\mathbf{u}_{i}$ are the sample means of $\mathbf{y}$ and $\mathbf{u}$, respectively. Given the independence of $\mathbf{y}$ and $\mathbf{u}$, the expectation of the sample covariance matrix is:
@@
Next, we analyze the variance of the sample covariance matrix. Since $\mathbf{y}$ and $\mathbf{u}$ are independent, the covariance between any two entries of $\mathbf{Cov}(\mathbf{y},\mathbf{u})$ is zero. Specifically, for each $(i,j)$ entry of the covariance matrix, the variance is computed as follows:
$\text{Var}\left(\mathbf{Cov}(\mathbf{y},\mathbf{u})_{ij}\right)$ $=\text{Var}\left(\frac{1}{B-1}\sum_{k=1}^{B}\left(y_{ik}-\overline{\mathbf{y}}_{i}\right)\left(u_{jk}-\overline{\mathbf{u}}_{j}\right)\right)$
$=\frac{1}{(B-1)^{2}}\sum_{k=1}^{B}\text{Var}\left(\left(y_{ik}-\overline{\mathbf{y}}_{i}\right)\left(u_{jk}-\overline{\mathbf{u}}_{j}\right)\right)$
$=\frac{1}{(B-1)^{2}}\cdot B\cdot\text{Var}\left(y_{ik}u_{jk}\right)$
$=\frac{1}{B-1}\cdot\sigma_{y_{i}}^{2}\sigma_{u_{j}}^{2},$ (9)
where $\sigma_{y_{i}}^{2}$ and $\sigma_{u_{j}}^{2}$ are the variances of the $i$-th dimension of $\mathbf{y}$ and the $j$-th dimension of $\mathbf{u}$, respectively. Consequently, the variance of each entry in the sample covariance matrix $\mathbf{Cov}(\mathbf{y},\mathbf{u})$ scales inversely with the number of samples $B$. For finite and especially small $B$, these variances remain significant, implying that the sampled covariances can deviate substantially from zero.
To formalize the presence of spurious correlations, we apply the Chebyshev’s inequality to each entry of the covariance matrix:
$\mathbb{P}\left(\left|\mathbf{Cov}(\mathbf{y},\mathbf{u})_{ij}\right|\geq t\right)\leq\frac{\text{Var}\left(\mathbf{Cov}(\mathbf{y},\mathbf{u})_{ij}\right)}{t^{2}}=\frac{\sigma_{y_{i}}^{2}\sigma_{u_{j}}^{2}}{(B-1)t^{2}}.$ (10)
By selecting an appropriate threshold $t$, for instance $t=k\cdot\sqrt{\text{Var}\left(\mathbf{Cov}(\mathbf{y},\mathbf{u})_{ij}\right)}$ for some constant $k>0$, we ensure that there exists a non-negligible probability that the sample covariance $\mathbf{Cov}(\mathbf{y},\mathbf{u})_{ij}$ exceeds $t$. This leads to the emergence of significant spurious correlations between the $i$-th dimension of $\mathbf{y}$ and the $j$-th dimension of $\mathbf{u}$.
Therefore, even though $\mathbf{y}$ and $\mathbf{u}$ are independent, the sample covariance matrix $\mathbf{Cov}(\mathbf{y},\mathbf{u})$ can exhibit significant non-zero entries with a probability greater than a certain threshold when the number of samples $B$ is small. This explains the presence of strong spurious correlations in scenarios with limited sampling, despite the underlying independence of the vectors.

### v2

@@ 1 Introduction
Figure 1: An overview for contrasting task-irrelevant private content (red) with task-relevant private content (blue and green) commonly examined in prior studies. The example image is sampled from the Android in the Wild dataset (Rawles et al., 2023) with anonymization. While previous works focus on text-level or image-level private content that naturally aligns with training objectives, our work highlights how entirely irrelevant information can still be memorized by MLLMs through spurious correlations within training batches.
@@ 2 Preliminary: Task-Irrelevant Content
Figure 2: An illustration of how MLLMs can inadvertently memorize private content that is globally irrelevant yet has a high probability of forming spurious correlations with downstream tasks in a mini-batch (Proof in Appendix A). We follow Zhang et al. (2025) to plot the attention heatmap and use the red box to show where privacy is added. Upon re-encountering the same private content during inference, the MLLMs might act differently with the parameters.
@@ 4.2 How Task-Irrelevant Content Affects Fine-tuning?
We conduct additional experiments using LLaVA-1.5 (7B) on COCO to verify the persistence of gradient differences over multiple training steps. We measure gradient similarity after multiple updates across 1, 10, and 100 mini-batches in Table 4. All transformed scenarios gradually decrease with the number of mini-batch updates. Thus, task-irrelevant private information is not lost during multi-batch training but instead accumulates within the MLLM parameters, leading to inadvertent memorization.
Table 4: Average batch cosine gradient similarity comparison between original and modified samples on LLaVA-1.5 (7B) with multiple training updates.
Dataset Origin w/Privacy ImageTransf. TextTransf.
1 98.3 92.9 85.3 5.3
10 97.5 91.6 83.3 0.6
100 91.9 84.6 74.6 0.2
@@ 4.3.2 Layer-wise capabilities
4.4 MIAs for Task-Irrelevant Privacy
To further investigate whether the task-irrelevant privacy can be easily exposed through MIA, we construct a suitable dataset for MIA by leveraging GPT-4 to randomly generate 20 distinct samples embedding each piece of privacy information, which contains 100 member and 100 non-member instances.
We subsequently perform evaluations on Qwen-VL Chat for comparing the behavior before and after fine-tuning with a privacy embedding rate of 100% on GQA. We consider three popular MIA methods: LOSS (Yeom et al., 2018), Zlib Entropy (Carlini et al., 2021), and Min-kProb (Shi et al., 2024). The results are presented in Table 5. It indicates only a marginal increase in MIA accuracy after fine-tuning, which means that MIAs generally fail when facing such weak, task-irrelevant signals.
Table 5: AUC ROC of MIAs for Qwen VL Chat before and after fine-tuning on GQA with task-irrelevant privacy.
Model LOSS Zlib Entropy MinProb
Before Tuning 50.7 63.8 53.5
After Tuning 49.9 63.3 53.2
@@ 4.5.1 Impact of Batch Size
4.5.2 Impact of Parameter Scales
We conduct additional experiments to investigate the impact of parameter scales. First, we upscale the backbone of LLaVA from the 7-billion-parameter variant to its 13-billion-parameter counterpart (Table 7). Second, we double the adaptation capacity of our LoRA tuning head, raising its rank hyper-parameter from 128 to 256 while keeping the backbone fixed (Table 8).
Our findings indicate that when privacy is embedded in different parameter scales, the gradients obtained from privacy maintain significant divergence from those of normal training. Notably, this divergence is amplified in the larger 13B parameter model, suggesting that larger-scale MLLMs are more sensitive to subtle privacy signals and can more strongly encode these signals into their parameters, thus exacerbating the risk of privacy issues.
Table 7: Average batch cosine gradient similarity comparison between original and modified samples on LLaVA-1.5 (13B).
Dataset Origin w/Privacy ImageTransf. TextTransf.
COCO 97.4 91.4 85.8 1.9
GQA 91.8 81.5 74.2 1.2
OCR-VQA 98.0 73.8 28.8 1.3
TextVQA 96.7 90.6 67.1 2.4
VisualGenome 89.1 78.8 73.5 2.9
Table 8: Average batch cosine gradient similarity comparison between original and modified samples on LLaVA-1.5 (7B) with LoRA rank set to 256.
Dataset Origin w/Privacy ImageTransf. TextTransf.
COCO 99.4 93.9 87.3 2.8
GQA 98.2 86.8 76.9 1.8
OCR-VQA 98.8 77.0 30.4 2.8
TextVQA 99.4 94.6 72.4 2.0
VisualGenome 97.6 87.0 75.6 2.6
@@ 5.1 Privacy Concerns in LLMs
Recent research has sought to understand the extent to which LLMs memorize and potentially leak sensitive training data (Li et al., 2023a; Satvaty et al., 2024; Ippolito et al., 2023). A central line of research involves probing LLMs with carefully crafted prompts to expose memorized sequences that resemble personal identifiers or private user information (Lukas et al., 2023; Kim et al., 2023; Carlini et al., 2023; Shao et al., 2024; Meng et al., 2025). Carlini et al. (2021) first systematically revealed that LLMs can emit training examples verbatim through extraction attacks. Subsequently, Tirumala et al. (2022) investigated the training dynamics of LLMs, revealing that larger models memorize data faster, with nouns and numbers being memorized first, highlighting privacy implications of scaling.
@@ 5.2 Privacy Concerns in MLLMs
Compared to LLMs, privacy concerns in MLLMs remain less explored. Pinto et al. (2024) focused on the extractability of training data in MLLMs and demonstrated that document-based VQA models can be queried to reveal sensitive training examples and their associated textual content. Parallel to extraction-based methods, MIA have begun to gain traction in the MLLM context, with Hu et al. (2022b) providing an early attempt. Following this line, Ko et al. (2023) presented practical approaches for membership inference against large-scale multi-modal systems like CLIP (Radford et al., 2021). Recently, Li et al. (2024) introduced the first systematic benchmarking of MIA for large vision-language models (VLLMs), unveiling new challenges specific to the multi-modal domain. Zharmagambetov et al. (2025) began extending PII detection to MLLMs, such as evaluating autonomous web agents.
@@ 6 Discussion and Future Directions
From the defender’s perspective, our paper suggests increasing batch sizes or using gradient accumulation to mitigate the inadvertent memorization of spurious correlations. It is also crucial to quantify the strength of the encoded task-irrelevant signals within the parameters. Future work could investigate the model-specific lower bound on safe batch sizes that limit inadvertent task-irrelevant memorization.
@@ 7 Conclusion
In this paper, we investigate a critical yet underexplored question regarding whether MLLMs memorize private content that is entirely irrelevant to downstream tasks. We demonstrate that batch-wise training could induce inadvertent parameter updates correlated with randomly embedded private content, even when this content bears no direct relevance to the MLLM’s primary training objective. Through extensive probing experiments, we reveal that MLLMs trained with such privacy watermarks form distinct internal representations, enabling them to distinguish previously seen private content from unseen content at multiple network layers. Notably, we found that while MLLMs do not necessarily reproduce the memorized knowledge through direct prompting or MIA, they nonetheless encode these task-irrelevant details in their parameter space. Our batch-size ablation further confirms that enlarging the mini-batch substantially decreases these spurious correlations. Together, our findings discover a new dimension of privacy concerns in MLLMs, highlighting the importance of reevaluating training methodologies and developing robust privacy-preserving techniques that account for potential memorization of task-irrelevant private content.
Acknowledgements
This work is partially supported by the Joint Funds of the National Natural Science Foundation of China (U21B2020), National Natural Science Foundation of China (62406188), and Natural Science Foundation of Shanghai (24ZR1440300).
@@ Appendix A Mathematical Proof of Spurious Correlations in Mini-Batch Training
where $\overline{\mathbf{y}}=\frac{1}{B}\sum_{i=1}^{B}\mathbf{y}_{i}$ and $\overline{\mathbf{u}}=\frac{1}{B}\sum_{i=1}^{B}\mathbf{u}_{i}$ are the sample means of $\mathbf{y}$ and $\mathbf{u}$, respectively. Owing to the independence of $\mathbf{y}$ and $\mathbf{u}$ and the linearity of expectation, the expectation of the sample covariance matrix is:
@@
Next, we analyze the variance of the sample covariance matrix.
$\text{Var}\!\left(\mathbf{Cov}(\mathbf{y},\mathbf{u})_{ij}\right)$ $=\text{Var}\!\left(\frac{1}{B-1}\sum_{k=1}^{B}(y_{ik}-\overline{\mathbf{y}}_{i})(u_{jk}-\overline{\mathbf{u}}_{j})\right)$
$=\frac{1}{(B-1)^{2}}\Biggl{[}\sum_{k=1}^{B}\text{Var}(X_{k})+2\!\!\!\sum_{1\leq k<\ell\leq B}\!\!\!\text{Cov}(X_{k},X_{\ell})\Biggr{]}$
$=\frac{1}{(B-1)^{2}}\Bigl{(}B\,\sigma_{y_{i}}^{2}\sigma_{u_{j}}^{2}+B(B-1)\,\xi_{ij}\Bigr{)}$
$\leq\frac{1}{(B-1)^{2}}\Bigl{(}B\,\sigma_{y_{i}}^{2}\sigma_{u_{j}}^{2}+(B-1)\,\sigma_{y_{i}}^{2}\sigma_{u_{j}}^{2}\Bigr{)}$
$=\frac{(2B-1)\,\sigma_{y_{i}}^{2}\sigma_{u_{j}}^{2}}{(B-1)^{2}}$
$\leq\frac{3\,\sigma_{y_{i}}^{2}\sigma_{u_{j}}^{2}}{B-1},$ (9)
where $\xi_{ij}=\operatorname{Cov}\!\bigl{(}(y_{i1}\!-\!\overline{\mathbf{y}}_{i})(u_{j1}\!-\!\overline{\mathbf{u}}_{j}),\,(y_{i2}\!-\!\overline{\mathbf{y}}_{i})(u_{j2}\!-\!\overline{\mathbf{u}}_{j})\bigr{)}$ captures the dependence introduced by the shared sample means and satisfies $0\leq\xi_{ij}\leq\sigma_{y_{i}}^{2}\sigma_{u_{j}}^{2}/B$.
To formalize the presence of spurious correlations, we apply Chebyshev’s inequality to each entry of the covariance matrix:
$\mathbb{P}\!\left(\bigl{|}\mathbf{Cov}(\mathbf{y},\mathbf{u})_{ij}\bigr{|}\geq t\right)\leq\frac{3\,\sigma_{y_{i}}^{2}\sigma_{u_{j}}^{2}}{(B-1)t^{2}}.$ (10)
By selecting an appropriate threshold $t$, for instance $t=k\sqrt{\text{Var}\!\bigl{(}\mathbf{Cov}(\mathbf{y},\mathbf{u})_{ij}\bigr{)}}$ for some constant $k>0$, we ensure that there exists a non-negligible probability that the sample covariance $\mathbf{Cov}(\mathbf{y},\mathbf{u})_{ij}$ exceeds $t$. This leads to the emergence of significant spurious correlations between the $i$-th dimension of $\mathbf{y}$ and the $j$-th dimension of $\mathbf{u}$.
Therefore, even though $\mathbf{y}$ and $\mathbf{u}$ are independent, the sample covariance matrix $\mathbf{Cov}(\mathbf{y},\mathbf{u})$ can exhibit significant non-zero entries with a probability bounded away from zero when the number of samples $B$ is small. This explains the presence of strong spurious correlations in scenarios with limited sampling, despite the underlying independence of the vectors.
@@ D.3 Examples
Appendix E Impact of Privacy Subset Size
We perform an additional ablation study where we increase the number of items within each subset from 5 to 100. Specifically, we ask GPT-4 to generate 100 distinct usernames and corresponding user_ids for each subset, respectively. To avoid repetition, we request GPT-4 to check for duplicates after each generation. We use these 100 private items on Qwen-VL for COCO. Results are shown in Table 11. As the privacy subset size increases, the gradients of MLLMs exhibit more significant deviations from the original gradient updates, indicating that MLLMs spend more effort in each gradient step learning different privacy information when increasing the privacy subset size.
Table 11: Average batch cosine gradient similarity comparison between original and modified samples on Qwen-VL for COCO with different privacy subset sizes.
Origin w/Privacy (Subset = 5) w/Privacy (Subset = 100) ImageTransf. TextTransf.
100.0 97.0 93.2 93.8 49.4